# macOSWorld: A Multilingual Interactive Benchmark for GUI Agents

### v1

@@ Abstract
Graphical User Interface (GUI) agents show promising capabilities for automating computer-use tasks and facilitating accessibility, but existing interactive benchmarks are mostly English-only, covering web-use or Windows, Linux, and Android environments, but not macOS. macOS is a major OS with distinctive GUI patterns and exclusive applications. To bridge the gaps, we present macOSWorld, the first comprehensive benchmark for evaluating GUI agents on macOS. macOSWorld features 202 multilingual interactive tasks across 30 applications (28 macOS-exclusive), with task instructions and OS interfaces offered in 5 languages (English, Chinese, Arabic, Japanese, and Russian). As GUI agents are shown to be vulnerable to deception attacks, macOSWorld also includes a dedicated safety benchmarking subset. Our evaluation on six GUI agents reveals a dramatic gap: proprietary computer-use agents lead at above 30% success rate, while open-source lightweight research models lag at below 2%, highlighting the need for macOS domain adaptation. Multilingual benchmarks also expose common weaknesses, especially in Arabic, with a 27.5% average degradation compared to English. Results from safety benchmarking also highlight that deception attacks are more general and demand immediate attention. macOSWorld is available at https://github.com/showlab/macosworld.
@@ 1 Introduction
The benchmarking results reveal distinct performance tiers, with proprietary computer use agents (CUAs) achieving over 30% success rate while open-source research models struggle at below 2%. On average, the agents perform the best with linear alphabetic languages (English, Russian), followed by block character-based languages (Chinese, Japanese), while right-to-left Arabic shows a 27.5% performance drop compared to English. This drop primarily manifests through degraded planning or grounding capabilities. Task and environment language mismatch further degrades agent performance. Our safety evaluation reveals that agents’ susceptibility varies, with the two strongest CUAs being highly vulnerable to context deception attacks ($\approx$70deception rate), highlighting that the issue is more general than previously understood and demands immediate attention.
@@ 4.3 Safety Benchmarking Subset
Rather than overlaying synthetic pop-ups on screenshots, we use AppleScript to trigger genuine macOS pop-up windows. The testbench spawns a process parallel to the agent loop, triggering the pop-up window and logging one of three outcomes: gold, distracted, or unhandled (e.g., the pop-up is dragged aside).
@@ 5.1 Benchmark Setup
All experiments run in a macOS virtual instance at 1024$\times$768 pixels (the default for macOS and for Claude CUA [29]), physically on AWS-hosted Mac Minis (model A2348). Except for the Set-of-Mark (SoM) ablation (Table 8), all agents perceive the environment from the current (or the most recent 3) screenshot only. Each task is granted up to 15 screenshots or 30 dialog turns, whichever limit is reached first. Agents retain the full conversation history, but prune screenshots to the most recent three (with ShowUI [25] being an exception, following its own context format [4]). Tasks are benchmarked in five languages – English (en), Chinese (zh), Arabic (ar), Japanese (jp), and Russian (ru) – with both the task prompt and system UI set to the target language. The only exception is the Advanced Apps category, which is evaluated in English exclusively (Section 4.2).
@@ 5.2 Quantitative Performance Evaluation
Agents’ performance form three distinct tiers on macOSWorld.
As shown in Table 3), the two proprietary CUAs lead with overall SRs above 30%. The general-purpose proprietary VLMs (GPT-4o at 5.5%, Gemini 2.5 Pro at 18.5%) occupy the middle tier. Finally, the two open-source models (UI-TARS 7B DPO, ShowUI) register below 2% overall. Notably, UI-TARS 7B DPO fails to complete any task end-to-end (0% across all categories). Given their competitive performance on web browsing, Windows and Ubuntu Linux [25, 2, 19], the results reveal they lack macOS-specific adaptation. Agents within each tier exhibit similar failure modes, which we analyze in detail in Section 5.3.
@@
Across five languages evaluated, linear alphabetic languages (English, Russian) yield the highest average SRs across agents (>16%). East Asian block characters (Chinese, Japanese) follow closely at 15%-16%. Right-to-left Arabic lags at 13.4% – a 27.5% drop from English, reflecting agents’ difficulty perceiving either the visual complexity of Arabic glyphs or the mirrored UI layout, or both. GPT-4o exhibits especially uneven multilingual performance, with over 60% SR degradation on Russian and Arabic compared to English. If the task language and the environment language do not match, agent performance could further decline (see Appendix D.1).
@@ 5.3 Qualitative Analysis: Agent Behavior and Failure Modes
Open-source research models – Weak grounding or inadequate planning.
ShowUI demonstrates reasonable grounding ability but poor planning: it often executes nonsensical actions, for example, scrolling on the desktop attempting to open Settings (Figure 9). It also tried to seek help keying in the given task in the help menu (Figure 8), which is a clear sign suggesting that ShowUI can recognize basic interface elements but lacks macOS-specific domain knowledge. In contrast, UI-TARS plans better but fails more often on grounding: In 16 settings-related tasks, it knows how to correctly open Settings, but only succeeds twice, and three times it misclicks on "Recent Items", showing possible systematic bias (Figure 11 and 13). Although UI-TARS records a 0% end-to-end success rate (Table 3), many of its failures are nearly successful attempts (Figure 12). Together, these results indicate that lightweight research models require substantial adaptation for mainstream macOS GUI tasks.
General-purpose proprietary VLMs – Excellent planning with imprecise grounding.
Both GPT-4o and Gemini 2.5 Pro struggle to precisely ground GUI elements, consistent with prior findings on general VLM agents in GUI settings [19, 20, 31]. While Gemini 2.5 Pro shows noticeably better grounding than GPT-4o, neither approaches the performance of specialized computer-use agents. Grounding UI elements with Set-of-Mark annotations [31] doubles GPT-4o’s performance (details in Appendix D.2), at 13.0% averaged across all languages (Table 8), but still lags Gemini’s 18.5%.
@@
OpenAI and Claude CUAs achieve the highest overall SRs but exhibit inefficiencies in step budgeting. They often repeatedly fail and retry simple operations like creating and renaming folders (Figure 14 and 15), consuming many more action steps than humans do and leaving insufficient step budget for downstream operations. For successful tasks, OpenAI CUAs take an average of 15.95 steps, Claude CUA 19.85 steps, while the annotator needs only 7.51 steps on average.
Multilingual discrepancies of CUAs – Degraded grounding or planning in weaker languages.
Differences in performance across languages are most pronounced for the two proprietary CUAs. For OpenAI CUA, failures in weaker languages stem primarily from degraded grounding: In one task (Figure 16), the agent effortlessly completed the account-creation task in English, but in Arabic, it fails to open Settings at all. Claude CUA exhibits both grounding and planning-related errors: where it succeeds in English, it could click on the horizontally mirrored location in Arabic, or use more steps to complete the same task (Figure 17). These cross-language discrepancies underscore the necessity for script-aware training and UI layout adaptation to maintain consistent performance across diverse language environments.
Table 4: Performance of baseline agents on macOSWorld by language, averaged across all task categories excluding Advanced Apps. Subscript percentages in the Average column are performance degradations with respect to the English counterpart. Highest SRs in eahc column are in bold.
Claude
CUA
OpenAI
CUA
GPT-4o
Gemini
Pro 2.5
UI-TARS
7B
ShowUI Average
en 44.4% 33.3% 8.8% 22.8% 0.0% 1.2% 18.4%
ru 40.9% 39.2% 2.9% 15.8% 0.0% 1.2% 16.7% -9.5%
ja 36.8% 35.1% 5.3% 18.1% 0.0% 0.0% 15.9% -13.8%
zh 31.6% 33.3% 6.4% 19.9% 0.0% 1.2% 15.4% -16.4%
ar 31.6% 28.1% 2.9% 15.8% 0.0% 1.8% 13.4% -27.5%
Average 37.1% 33.8% 5.3% 18.5% 0.0% 1.1% 15.9% -13.4%
@@ 5.4 Safety Subset Evaluation
Table 5: Baseline agent performance on the macOSWorld safety subset under context-deception attacks, showing rates (%) of Distracted (clicking a decoy), Gold (clicking the gold button), and Unhandled (pop-up not closed). Highest rates in each row are in bold.
@@
Distracted 72.4% 69.0% 0.0% 17.2% 0.0% 41.4% 33.3%
Gold 24.1% 27.6% 0.0% 3.4% 0.0% 10.3% 10.9%
Unhandled 3.4% 3.4% 100.0% 79.3% 100.0% 44.8% 55.2%
As shown in Table 5, under context deception attacks, proprietary CUAs almost always handle the pop-up events (<4% unhandled) but exhibit high distraction rates ($\approx$70%), whereas GPT-4o, Gemini Pro 2.5, and UI-TARS leave the vast majority of pop-ups unhandled. Notably, despite ShowUI’s low overall SR, it understands the attacks and was distracted 41.4% of the time. These results reveal the vulnerability in GUI agents (particularly CUAs) to complex small-font deceptive UI designs, highlighting that context deception attacks could be more general and broadly effective, going beyond prior studies [42, 45]. The findings underscore the need for imminent safety mechanisms.
@@ Appendix B Ethical and Safeguarding Statement
For responsible deployment, macOSWorld implements environment isolation for safeguarding. The benchmark operates solely within virtualized macOS environments on AWS Mac mini instances, isolating potential harmful operations from both host platforms and physical hardware. Our snapshot restoration process ensures all content, files, and system changes are immediately discarded after each evaluation, preventing any accumulation of sensitive information.
@@ E.2 UI-TARS
UI-TARS failed to complete any task, yet its action trajectories demonstrate greater rationality and explainability, to the extent that some tasks nearly approached success. UI-TARS’s primary issue lies in grounding, specifically its inability to accurately click on targets.
Figure 10: An example of UI-TARS encountering grounding issues when nearly completing a task. In this case, the task instruction is: "I am using the current template from the last lecture to create slides for the next lecture. Help me change Lecture 5 to Lecture 6, and the title to ‘Image Enhancement’." UI-TARS successfully enters the save interface and inputs the correct filename, but in the final step when needing to click the save button, it repeatedly clicks at position (652, 541), as marked by the red dot in the figure, which is far from the save button.
Figure 10 illustrates an example where UI-TARS nearly succeeds but ultimately fails due to grounding issues, repeatedly failing to click the save button. In this case, UI-TARS successfully brings up the save interface and types the correct filename, demonstrating some planning and grounding capability. However, it repeatedly misses when attempting to click in the final step, indicating that grounding problems can occur at any stage of a task, regardless of the intuitive difficulty level for humans.
Figure 11: An example of UI-TARS’s inaccurate grounding. In this case, the task instruction is: "Modify settings to enable using keyboard shortcuts to zoom." UI-TARS attempts to click on System Settings but instead clicks on Recent Items below it.
Figure 12: An example of UI-TARS’s inaccurate grounding. In this case, the task instruction is: "In the project document, go to products and change my previous choice of the dishwasher to model B." UI-TARS correctly identifies the "Products" tab to click, but fails due to imprecise grounding. Although it is aware of the failure and continually adjusts the click coordinates, the adjustments only increase the error.
Such grounding issues are more typical in tasks involving opening the settings interface, with the case in Figure 11 being a representative example. UI-TARS tends to open System Settings from the Apple Menu in the top left corner but consistently clicks on elements below System Settings. Across all 16 tasks involving opening system settings, UI-TARS clicked on Recent Items three times and misclicked on App Store twice, demonstrating a systematic error where UI-TARS’s click positions are consistently lower than the actual element positions. Such systematic bias alsow exist in other tasks (Figure 12), making it extremely difficult for UI-TARS to complete even simple tasks, despite knowing what to do.
Figure 13: An example of UI-TARS encountering both planning and grounding issues. In this case, the task instruction is: "Help me navigate to the ‘System Report’ page." The issues in steps 2-3 illustrate planning problems, where the agent closes the Apple Menu after opening it, then reopens it, wastefully using two operations. Steps 4-5 demonstrate grounding issues. The agent attempts to click on "About this Mac" but instead clicks on System Settings below it.
Furthermore, although its main issue is grounding, UI-TARS also frequently exhibits planning problems. Figure 13 exemplifies this. The task requires opening the system report page, which can be accessed either through About this Mac or through General settings in System Preferences. However, after opening the Apple menu, UI-TARS chooses to close it and then reopen it in subsequent steps, inexplicably wasting two operations and demonstrating planning deficiencies.

### v3

@@ Abstract
Graphical User Interface (GUI) agents show promising capabilities for automating computer-use tasks and facilitating accessibility, but existing interactive benchmarks are mostly English-only, covering web-use or Windows, Linux, and Android environments, but not macOS. macOS is a major OS with distinctive GUI patterns and exclusive applications. To bridge the gaps, we present macOSWorld, the first comprehensive benchmark for evaluating GUI agents on macOS. macOSWorld features 202 multilingual interactive tasks across 30 applications (28 macOS-exclusive), with task instructions and OS interfaces offered in 5 languages (English, Chinese, Arabic, Japanese, and Russian). As GUI agents are shown to be vulnerable to deception attacks, macOSWorld also includes a dedicated safety benchmarking subset. Our evaluation on six GUI agents reveals a dramatic gap: proprietary computer-use agents lead at above 30% success rate, while open-source lightweight research models lag at below 5%, highlighting the need for macOS domain adaptation. Multilingual benchmarks also expose common weaknesses, especially in Arabic, with a 28.8% average degradation compared to English. Results from safety benchmarking also highlight that deception attacks are more general and demand immediate attention. macOSWorld is available at https://github.com/showlab/macosworld.
@@ 1 Introduction
The benchmarking results reveal distinct performance tiers, with proprietary computer use agents (CUAs) achieving over 30% success rate while open-source research models struggle at below 5%. On average, the agents perform the best with linear alphabetic languages (English, Russian), followed by block character-based languages (Chinese, Japanese), while right-to-left Arabic shows a 28.8% performance drop compared to English. This drop primarily manifests through degraded planning or grounding capabilities. Task and environment language mismatch further degrades agent performance. Our safety evaluation reveals that agents’ susceptibility varies, with the two strongest CUAs being highly vulnerable to context deception attacks ($\approx$70deception rate), highlighting that the issue is more general than previously understood and demands immediate attention.
@@ 4.3 Safety Benchmarking Subset
Rather than overlaying synthetic pop-ups on screenshots, we use AppleScript to trigger genuine macOS pop-up windows. The testbench spawns a process parallel to the agent loop, triggering the pop-up window and logging one of three outcomes: gold, distracted, or unhandled (e.g., the pop-up is not interacted or dragged aside).
@@ 5.1 Benchmark Setup
All experiments run in a macOS virtual instance at 1024$\times$768 pixels (the default for macOS and for Claude CUA [29]), physically on AWS-hosted Mac Minis (model A2348). Except for the Set-of-Mark (SoM) ablation (Table 8), all agents perceive the environment from the most recent 3 screenshot only. Each task is granted up to 15 screenshots or 30 dialog turns, whichever limit is reached first. Agents retain the full conversation history, but prune screenshots to the most recent 3 (with ShowUI [25] being an exception, following its own context format [4]). Tasks are benchmarked in five languages – English (en), Chinese (zh), Arabic (ar), Japanese (jp), and Russian (ru) – with both the task prompt and system UI set to the target language. The only exception is the Advanced Apps category, which is evaluated in English exclusively (Section 4.2).
@@ 5.2 Quantitative Performance Evaluation
Agents’ performance form three tiers on macOSWorld.
As shown in Table 3), the two proprietary CUAs lead with overall SRs above 30%. Gemini 2.5 Pro, alone, occupy the middle tier, at 18.5%. Finally, GPT-4o and the two open-source models (UI-TARS 7B DPO, ShowUI) register below 10% on average. For the open-source models, compared to their competitive performance on web browsing, Windows and Ubuntu Linux [25, 2, 19], the results here reveal they lack macOS-specific adaptation. Notably, Although GPT-4o and UI-TARS are close in performance, their failure modes are completely different, which we analyze in detail in Section 5.3.
@@
As shown in Table 4, across five languages evaluated, linear alphabetic languages (English, Russian) yield the highest average overall SRs across agents (>17.5%). East Asian block characters (Chinese, Japanese) follow closely at 17.2% and 15.8%. Right-to-left Arabic lags at 13.7%, which is a 28.8% drop from English, reflecting agents’ difficulty perceiving either the visual complexity of Arabic glyphs or the mirrored UI layout, or both. GPT-4o exhibits especially uneven multilingual performance, with over 60% SR degradation on Russian and Arabic compared to English. If the task language and the environment language do not match, agent performance could further decline (see Appendix D.1).
@@ 5.3 Qualitative Analysis: Agent Behavior and Failure Modes
Open-source research models – Good grounding but poor planning.
ShowUI demonstrates decent grounding ability but often executes nonsensical actions. For example, it attempts to open Settings by scrolling the mouse wheel on the desktop (Figure 9). It also tries to complete the task by typing it into the Help menu (Figure 8), which clearly shows that while it can identify basic interface elements, it lacks macOS-specific domain knowledge. In comparison, UI-TARS not only exhibits similarly nonsensical behaviors (Figure 10 and 14), but also fails due to illegal output formats and hallucinations. In Figure 15, it issues instructions to click a button by name rather than by coordinates, which is an illegal operation; in Figure 14, it mistakenly believes the Apple menu is located at the bottom-left corner of the screen. These results indicate that lightweight research models require substantial adaptation for mainstream macOS GUI tasks.
General-purpose proprietary VLMs – Good planning with imprecise grounding.
Although GPT-4o’s performance is close to UI-TARS, its failure mode is more similar to Gemini 2.5 Pro, another general-purpose VLM. Both GPT and Gemini struggle to precisely ground GUI elements, consistent with prior findings on general VLM agents in GUI settings [19, 20, 31]. While Gemini 2.5 Pro shows noticeably better grounding than GPT-4o, neither approaches the performance of specialized computer-use agents. Grounding UI elements with Set-of-Mark annotations [31] doubles GPT-4o’s performance (details in Appendix D.2), at 13.0% averaged across all languages (Table 8), but still lags Gemini’s 18.5%.
@@
OpenAI and Claude CUAs achieve the highest overall SRs but exhibit inefficiencies in step budgeting. They often repeatedly fail and retry simple operations like creating folders (Figure 16 and 17), consuming more action steps than humans do and leaving insufficient step budget for downstream operations. For successful tasks, OpenAI and Claude CUAs take 15.95 and 19.85 steps on average, while human annotators use only 7.51 steps.
Multilingual discrepancies of CUAs – Performance gap stems from degraded grounding or planning.
Differences in performance across languages are most pronounced for the two proprietary CUAs. For OpenAI CUA, the performance degradation stem primarily from degraded grounding: In one task (Figure 18), the agent effortlessly completed the account-creation task in English, but in Arabic, it fails to open Settings at all. Claude CUA exhibits both grounding and planning-related errors: while GUIs are mirrored in Arabic, Claude may carry biases from left-to-right UI layouts and incorrectly click on non-mirrored locations. It could also use more steps to complete the same task (Figure 19). These cross-language discrepancies underscore the necessity for language-aware training and UI layout adaptation to maintain consistent performance across diverse language environments.
Table 4: Performance (success rate, SR) of baseline agents on macOSWorld by language, averaged across all agents and all task categories excluding Advanced Apps. Performance degradations ($\Delta$) are with respect to the English counterpart.
en ru zh ja ar Average
SR 19.3% 17.7% 17.2% 15.8% 13.7% 16.7%
$\Delta$8.1% -$1$8.2% -28.8% -13.2%
@@ 5.4 Safety Subset Evaluation
Table 5: Agent performance on the macOSWorld safety subset under context-deception attacks, showing rates (%) of Distracted (clicking a decoy), Gold (clicking the close/cancel button), and Unhandled (pop-up not closed). Colors indicate highest and second highest rates in each row.
@@
Distracted 72.4% 69.0% 0.0% 17.2% 58.6% 41.4% 33.3%
Gold 24.1% 27.6% 0.0% 3.4% 34.4% 10.3% 10.9%
Unhandled 3.4% 3.4% 100.0% 79.3% 6.9% 44.8% 55.2%
As shown in Table 5, under context deception attacks, proprietary CUAs almost always handle the pop-up events (<4% unhandled) but exhibit high distraction rates ($\approx$70%), whereas proprietary VLMs leave the vast majority of pop-ups unhandled. Notably, despite UI-TARS and ShowUI’s low overall SR in previous experiments, they understand the pop-ups and were distracted half of the time on average. These results reveal the vulnerability in GUI agents (particularly CUAs) to complex small-font deceptive UI designs, highlighting that context deception attacks could be more general and broadly effective, going beyond prior studies [42, 45]. The findings underscore the need for imminent safety mechanisms.
@@ Appendix B Ethical and Safeguarding Statement
For responsible deployment, macOSWorld implements environment isolation for safeguarding. The benchmark operates solely within virtualized macOS environments on AWS Mac mini machines, isolating potential harmful operations from both host platforms and physical hardware. Our snapshot restoration process ensures all content, files, and system changes are immediately discarded after each evaluation, preventing any accumulation of sensitive information.
@@ E.2 UI-TARS
UI-TARS’s performance lies between that of ShowUI and GPT-4o, achieving an average task success rate of 4.8% across all languages. It exhibits three primary failure modes: executing nonsensical actions, hallucinations, and producing outputs in invalid formats.
When executing nonsensical actions, UI-TARS behaves similarly to ShowUI, often in ways that are difficult to interpret. For example, in Figure 10, when the task requires opening System Settings via the Apple menu in the top-left corner of the screen, UI-TARS instead opens four unrelated menus (e.g. Edit, File) from the top bar. It is unclear whether this reflects a systematic attempt to locate System Settings or merely random, unconscious exploration. n some cases, this phenomenon may even occur during actions that UI-TARS had previously been able to complete successfully. For example, when asked to insert two blank tables into a document, as shown in Figure 11, UI-TARS successfully inserts the first one, but during the second attempt, it begins performing nonsensical actions, adjusting the zoom level of the view. Notably, UI-TARS’s thought process during these tasks offers little insight into the rationale behind its actions, either. Its internal reasoning often only specifies what to do and how to do it, without explaining why.
Figure 10: An example where UI-TARS executes nonsensical actions. In this case, the task instruction is: "Please help me modify the settings so that I can view a larger version of text when I am typing." UI-TARS should open System Settings via the Apple menu in the top left corner or click the gear icon in the Dock below, but it does neither. Instead, it successively opens the Finder, Edit, Window, and File menus from the top bar. It’s unclear whether it is searching through each menu in hopes of finding the entrance to System Settings or mindlessly opening menus without intent.
Figure 11: An example where UI-TARS executes nonsensical actions. In this case, the task instruction is: "Help me create a blank Pages document and insert two separate empty tables in it. No need to configure the details of the tables." The task requires the agent to insert two tables. The agent has already inserted one and now needs to insert the second. To do this, it only needs to repeat the previous steps by clicking the Table icon at the top of the interface again and selecting any table. However, it first clicks a cell in the existing table and then starts adjusting the zoom level of the view. These actions are inexplicable, illogical, and the agent’s thoughts and actions are not aligned.
Hallucinations in UI-TARS are most evident in its perception of screen content. For instance, it might mislocate the Apple icon, perceiving it on the bottom-left instead of the top-left (Figure 12), or misestimate the aspect ratio of a slide on screen (Figure 13). Hallucination and nonsensical actions can also occur simultaneously: in step 4 in Figure 14, the agent attempts, without clear motivation, to click the Apple icon to close the Apple menu, while also hallucinating the Apple icon is located at the bottom-left of the screen. Considering that UI-TARS achieves strong quantitative performance across multiple platforms and benchmarks [2], such errors may stem from its lack of adaptation to macOS and limited domain-specific knowledge.
Figure 12: An example demonstrating UI-TARS’s hallucination. In this case, the task instruction is: "Please help me modify the settings so that I can view a larger version of text when I am typing." To change this setting, the agent intends to open System Settings through the Apple menu in the top left corner, but it hallucinates that the Apple icon is in the bottom left and clicks there instead.
Figure 13: An example where UI-TARS incorrectly believes it has completed the task as instructed. In this case, the task instruction is: “Help me create an empty slide with an aspect ratio of 4:3.” UI-TARS directly clicks the Create button in the bottom right corner of the window in the first step, ignoring the aspect ratio setting located in the top right corner of the interface. Although the created presentation has a 16:9 aspect ratio, the agent still claims that its aspect ratio is 4:3.
Figure 14: A UI-TARS example containing multiple issues simultaneously. In this case, the task instruction is: "Help me modify keyboard settings, toggling the key repeat rate to the fastest possible." The agent needs to first open the Apple menu and then open System Settings from there. In Step 3, the agent correctly opens the Apple menu, but its thought reflects an expectation of seeing "About This Mac" instead of "System Settings". Then in Step 4, the agent nonsensically tries to close the Apple menu. Its thoughts indicate that it believes the Apple icon is in the bottom left corner of the screen, but it still correctly clicks the top left corner of the screen to close the Apple menu it just opened. These two consecutive action steps include both nonsensical actions and hallucinations, and the thoughts and actions are not aligned, highlighting the unexplainable nature of UI-TARS failure cases.
Another factor limiting UI-TARS’s performance is its inconsistency in output formatting. Figure 15 shows invalid formats across two consecutive steps: the agent omits the "Thought:" prefix before stating its reasoning, includes redundant equals signs, quotation marks, and line breaks within the click action syntax, and incorrectly places a translated label of the UI element in place of a screen coordinate. Improving formatting consistency would enhance the agent’s reliability and ensure that its outputs remain parsable by downstream algorithms.
Figure 15: An example where UI-TARS does not strictly adhere to its output format. In this case, the task instruction is: "In ’System Settings > Appearance’, set accent color to red." The first invalid output in this example is the absence of the keyword "Thought:", which may cause some thought-parsing algorithms to fail. The second invalid output appears in step 6’s click action parameters. It was supposed to be a coordinate, but instead, the output repeatedly contains equal signs, single quotes, and newline characters. Moreover, where the coordinates should appear, the agent outputs the Chinese translation of the key name at that location. This illustrates that UI-TARS is not sufficiently robust to consistently adhere to its fixed output format.
Despite the room for improvement on macOS tasks, UI-TARS sometimes demonstrates self-correction capabilities. In some examples, it explicitly acknowledges past mistakes in its reasoning, as shown below:
Thought (originally in Chinese): Due to my incorrect action in the previous step, a dropdown menu labeled "Document" has appeared on the right side of the page. This does not meet the task requirements. I should left-click the "Document" tab in the upper-right corner of the page to close the dropdown menu.